Make New Entry,
Make Followup Entry
User name Yi Qiang
Log entry time 10:18:07 on November 11, 2008
Entry number 248118
This entry is a followup to: 248113
The correction is based on the Lock-in output. I found out (after I
turned off the laser and oven :S ) that the lock-in output on the high
frequency side has a -150 kHz off set which means there is also an offset
in the frequency. And the slope we obtained from FM test is
1.028uV/kHz*10000mV/500uV = 20.6 mV/kHz, therefore the frequency offset
is -150/20.6 = -7.28 kHz.
The frequency difference in the EPR test is 47.73 corresponding to 38.7%
polarization. After taking this offset into account, the real frequency
difference is 47.73+7.28 = 55.01 and the polarization is 44.6%.
So the calibration constant is also corrected from 1.258 to 1.45 %/mV.
The registry has been updated with the corrected value.
A copy of this log entry has been emailed to: email@example.com, firstname.lastname@example.org, email@example.com, firstname.lastname@example.org